# CVOCSemRPL: Class-Variance Optimized Clustering, Semantic Information Injection and Restricted Pseudo Labeling based Improved Semi-Supervised Few-Shot Learning

### v2

@@ 4.3.2 Episodic fine-tuning losses
Once CVOC has delivered a set of refined prototypes and the corresponding query logits $Q_{\text{cvoc}}\!\in\!\mathbb{R}^{(N_{q}C)\times C}$, we optimise the visual extractor $M$ with three episode-level losses:
@@
$L_{\mathrm{cvoc}}=\operatorname{CE}(Q_{\text{cvoc}},y_{q}).$ (13)
In parallel, we pass the same query embeddings through the label propagation component  [24], producing logits $Q_{\text{lp}}$; the resulting loss is as follows.
$L_{\mathrm{lp}}=\operatorname{CE}(Q_{\text{lp}},y_{q}).$ (14)
@@
$L_{\mathrm{cls}}=\operatorname{CE}(Q_{\text{all}},y_{\text{all}}).$ (15)
@@ 4.4 Training Semantic Injection Network
Figure 2: (Left) Semantic Injection Network. (Right) Evaluation using semantic injection and restricted pseudo-labeling.
[…]
@@
Where, $F_{L1}$ refers to L1-loss, $P_{c_{j}}$ refers to the class $c_{j}$ prototype, $N_{c_{j}}$ refers to the number of labeled samples belonging to class $c_{j}$. We introduce a simple, automated pipeline for extracting high-quality, visually informative class descriptions. Please refer to the supplementary material for details.
@@ 4.5 Evaluation with Semantic Injection and Restricted Pseudo-Labeling
After pre-training and episodic fine-tuning, we evaluate on $N$-way $K$-shot test episodes composed of a labeled support set $\mathcal{S}_{\text{test}}$, an unlabeled pool $\mathcal{U}_{\text{test}}$, and a query set $\mathcal{Q}_{\text{test}}$. We first run CVOC on $\mathcal{S}_{\text{test}}\!\cup\!\mathcal{U}_{\text{test}}$ to obtain visual centres $P(y)$. For each class $y$ we encode its textual description with the CLIP text encoder and obtain a semantic vector $t_{y}$ [27]. Concatenating $P(y)$ with $t_{y}$ and processing the pair through the semantic-injection encoder $E_{s}$ yields the refined prototype
@@
where $s\in[0,1]$ is a fixed blend weight. This constitutes the overall procedure of our semantic anchor module — aligning visual prototypes via the semantic injection network and prototype correction according to Eq. (22).
For each $x_{i}^{u}\in\mathcal{U}_{\text{test}}$ we compute the entropy $H(x_{i}^{u})$ of its softmax logits and retain only the lowest-entropy $k\%$. The selected subset and its hard pseudo-labels are
@@
The support set is then expanded with these high-confidence examples:
$\mathcal{S}^{0}=\mathcal{S}_{\text{test}}\cup\widetilde{\mathcal{U}}_{\text{test}}.$ (24)
Balanced label propagation [24] is applied to the features of $\mathcal{S}^{0}\cup\mathcal{Q}_{\text{test}}$, using the labels in $\mathcal{S}^{0}$ as anchors. The propagated soft labels for $\mathcal{Q}_{\text{test}}$ are finally hardened by selecting the highest-probability class, yielding the episode predictions (Fig. 4).
4.6 Quality of Representation Space
Prior methods in SSFSL often commit to either global prototype-based structures (e.g., ProtoNet [10], Soft K-Means [28]) or reconstruction-based alignments (e.g., MFC [4]). Prototype-based methods, while offering clear global decision boundaries, can falter in capturing fine-grained intra-class variations and local geometric nuances. Conversely, reconstruction-based clustering excels at modelling local manifolds but often mis-assign samples when two class subspaces overlap, because they ignore where class centroids sit in the broader topology. We overcome this dichotomy by fusing the reconstruction based distance with explicit Euclidean regularisers: an intra-class term that contracts each cluster and an inter-class term that repels different centroids. A point is accepted by a class only if it is simultaneously well reconstructed by that class’s subspace and lies near its prototype while far from others. This fusion retains the fine-grained shape of each class manifold while simultaneously positioning the manifolds far enough apart to maintain clear global separation. In the fine-tuning stage, the cluster separation tuner (CST) adjusts each centroid to maximize intra-class compactness and inter-class separation.
Next, we nudge each visual prototype a small step toward the CLIP [27] text embedding of its class description using our semantic injection network. This language-based anchor acts like a compass: it steers prototypes away from visually misleading regions (e.g., background clutter or occlusion) and pulls them toward positions that reflect the class’s true meaning. That is particularly valuable when we have only a handful of labeled examples.
Collectively, the geometry-aligned fusion, semantic anchoring, and margin polishing generate a latent space that is tighter within classes, farther between classes, and conceptually aligned with the class description. This enhanced discriminability yields substantially higher pseudo-label accuracy than prior approaches (see Table 5, where we compared our method against Kmeans and MFC [4]), which in turn drives significant gains in FSL performance.
@@ 4.6 Quality of Representation Space
Method 1-shot 5-shot Method 1-shot 5-shot
TADAM 58.50$\pm$0.3076.70$\pm$0.30LEO 61.76$\pm$0.0877.59$\pm$0.12%
CAN 67.19$\pm$0.5580.64$\pm$0.35wDAE-GNN 62.96$\pm$0.6278.85$\pm$0.10%
LST 70.10$\pm$1.9078.70$\pm$0.80EPNet 80.74$\pm$0.8584.34$\pm$0.53%
EPNet 66.50$\pm$0.8981.06$\pm$0.60TransMatch 62.93$\pm$1.1182.24$\pm$0.59%
TPN 59.46% 75.65% ICI 71.41% 81.12%
PLAIN 74.38$\pm$2.0682.02$\pm$1.08EPNet-SSL 79.22$\pm$0.9288.05$\pm$0.51%
EPNet-SSL 75.36$\pm$1.0184.07$\pm$0.60PTN 81.57$\pm$0.9487.17$\pm$0.58%
cluster-FSL 77.81$\pm$0.8185.55$\pm$0.41cluster-FSL 82.63$\pm$0.7989.16$\pm$0.35%
CVOCSemRPL 84.51$\pm$0.5486.95$\pm$0.39CVOCSemRPL 85.94$\pm$0.5491.03$\pm$0.34%
Table 2: Classification accuracy (%) on the tieredImageNet dataset
@@
Method 1-shot 5-shot Method 1-shot 5-shot
MetaOpt-SVM 65.99$\pm$0.7281.56$\pm$0.53Soft K-Means 51.52$\pm$0.3670.25$\pm$0.31%
CAN 73.21$\pm$0.5884.93$\pm$0.38LEO 61.76$\pm$0.0878.43$\pm$0.12%
LST 77.70$\pm$1.6085.20$\pm$0.80wDAE-GNN 68.16$\pm$0.4682.96$\pm$0.51%
EPNet 76.53$\pm$0.8384.80$\pm$0.98EPNet 78.50$\pm$0.3188.90$\pm$0.57%
PLML-EP 79.62% 86.69% ICI 85.44% 89.12%
PLAIN 82.91$\pm$2.0989.28$\pm$1.25EPNet-SSL 83.68$\pm$0.6989.00$\pm$0.52%
EPNet-SSL 81.79$\pm$0.9788.45$\pm$0.61PTN 84.70$\pm$1.5489.14$\pm$0.73%
cluster-FSL 83.89$\pm$0.8189.94$\pm$0.46cluster-FSL 85.74$\pm$0.7690.18$\pm$0.43%
CVOCSemRPL 85.55$\pm$0.6591.13$\pm$0.31CVOCSemRPL 89.06$\pm$0.5891.53$\pm$0.32%
@@ 5 Experiments
The SSFSL methods are evaluated on the miniImageNet [29], tieredImageNet [28], and CUB [30] datasets. Please refer to the supplementary material for further details on the datasets used, the implementation details, and the compared methods.
@@ 5.1 Results
We conducted experiments under 5-way 1-shot and 5-way 5-shot settings, on the miniImageNet, tieredImageNet, and CUB-200-2011 datasets, using WRN-28-10 and ResNet-12 backbones. For the miniImageNet dataset (Table 1), our CVOCSemRPL model outperformed cluster-FSL [4], achieving a 6.7% and 1.4% improvement in 1-shot and 5-shot settings with ResNet-12, and a 3.31% and 1.87% improvement with WRN-28-10. On the tieredImageNet dataset (Table 2), CVOCSemRPL outperforms the closest method by margins of 1.66% and 1.19% for 1-shot and 5-shot cases with ResNet-12, and by a margin of 3.32% for 1-shot and 1.35% for 5-shot settings with WRN-28-10. On the CUB-200-2011 dataset (Table 3), CVOCSemRPL outperformed EPNet [3] and cluster-FSL [4] by significant margins. All experiments were conducted on 1000 few-shot tasks with 100 unlabeled data, and 95% confidence intervals are reported.
@@
Method 1-shot 5-shot 1-shot 5-shot
EPNet 82.85$\pm$0.8191.32$\pm$0.4187.75$\pm$0.7094.03$\pm$0.33%
cluster-FSL 87.36$\pm$0.7192.17$\pm$0.3191.80$\pm$0.5895.07$\pm$0.23%
CVOCSemRPL (ours) 88.63$\pm$0.6393.37$\pm$0.1593.15$\pm$0.4196.25$\pm$0.24%
@@
$\times$ $\times$ 78.31$\pm$0.7985.75$\pm$0.41%
$\checkmark$ $\times$ 80.29$\pm$0.5485.91$\pm$0.79%
$\times$ $\checkmark$ 82.53$\pm$0.5286.45$\pm$0.41%
$\checkmark$ $\checkmark$ 84.51$\pm$0.5486.95$\pm$0.39%
@@ 5.2 Ablation Experiments
We experimentally analyze the contributions of components of our method on the miniImageNet dataset and ResNet12 architecture. The results in Table 5 indicate that removing the semantic injection, or intra-inter class distance, significantly degrades the performance of our method. Fig. 3 indicates the utility of CST. Please refer to the supplementary material for other ablation experiments.
@@ 6 Conclusion
In this work, we proposed a novel SSFSL approach that performs improved clustering-based pseudo-labeling that promotes better intra/inter-class alignment and leverages semantic anchoring. Extensive experiments demonstrate that our method significantly outperforms recent state-of-the-art approaches. Limitations: Our approach assumes that both labeled and unlabeled data lie on a shared manifold, which is crucial for accurate pseudo-labeling. Our semantic anchoring technique relies heavily on the quality of raw class descriptions, which must provide rich visual characteristics to enable effective discrimination, and the choice of embedding model used to encode these descriptions. Consequently, this semantic anchor module may not work very well in complex datasets where the visual features of different classes are highly similar. This module also introduces some computational overhead.
Supplementary Material
7 Experimental Details
7.1 Datasets

### v3

@@ 4.3.2 Episodic fine-tuning losses
Once CVOC has delivered a set of refined prototypes and the corresponding query logits $Q_{\text{cvoc}}\!\in\!\mathbb{R}^{(N_{q}C)\times C}$, we optimize the visual extractor $M$ with three episode-level losses:
@@
$L_{\mathrm{cvoc}}=L_{\mathrm{ce}}(Q_{\text{cvoc}},y_{q}).$ (10)
In parallel, we pass the same query embeddings through the label propagation component [iscen2019label], producing logits $Q_{\text{lp}}$. The resulting loss is as follows.
$L_{\mathrm{lp}}=L_{\mathrm{ce}}(Q_{\text{lp}},y_{q}).$ (11)
@@
$L_{\mathrm{cls}}=L_{\mathrm{ce}}(Q_{\text{all}},y_{\text{all}}).$ (12)
@@ 4.4 Training Semantic Injection Network
Figure 2: Semantic Injection Network
Once the pre-training and finetuning of the visual feature extractor ($M$) are completed, we train a semantic injection network ($S$). For training this network, we need class descriptions. We propose a simple, automated class description generation pipeline to generate high-quality, visually informative class descriptions (refer to Sec. 8.4.4 in the Appendix). This is done to avoid any ambiguities that may arise from using just the class names. We use CLIP [radford2021learning] text encoder to obtain the class embeddings/semantic features from the generated class descriptions. We extract features of images $x_{i}$ from the train set using the fine-tuned visual feature extractor $M$ and embedding propagation. We obtain the class prototypes $P_{c_{j}}$ by simply averaging the features from corresponding training samples from each class $c_{j}$. We concatenate the class prototype with the semantic features of the same class and pass the concatenated it to the encoder module $E_{s}$ of the semantic injection network to obtain a joint embedding $E_{s}\bigl(P_{c_{j}},\,s_{c_{j}}\bigr)$. The decoder module $D_{s}$ of the semantic injection network is used to reconstruct the concatenated feature that was passed into $E_{s}$.
We train the semantic injection network S using three loss functions: joint embedding prototype loss $L^{jep}_{s}$, image feature prototype reconstruction loss $L^{fr}_{s}$, and semantic feature reconstruction loss $L^{sr}_{s}$ shown in Eq. 14. The joint embedding prototype loss $L^{jep}_{s}$ helps move the joint embedding closer to the image feature prototype of the same class (see Eq. 15). The image feature prototype of a class is computed by taking the mean of the features of images belonging to that class (see Eq. 16). The image feature prototype reconstruction loss $L^{fr}_{s}$ is used to reduce the difference between the reconstructed prototype $P^{r}_{c_{j}}$ and the actual prototype feature $P_{c_{j}}$. Similarly, the semantic feature reconstruction loss $L^{sr}_{s}$ is used to reduce the difference between the reconstructed semantic feature $s^{r}_{c_{j}}$ and the actual semantic feature $s_{c_{j}}$. The $L^{fr}_{s}$ and $L^{sr}_{s}$ losses force the semantic injection network to incorporate semantic information into the encoded prototype produced by $E_{s}$ (see Fig. 2).
@@
Where, $F_{L1}$ refers to L1-loss, $P_{c_{j}}$ refers to the class $c_{j}$ prototype, $N_{c_{j}}$ refers to the number of labeled samples belonging to class $c_{j}$.
Figure 3: Evaluation using semantic injection and restricted pseudo-labeling
@@ 4.5 Evaluation with Semantic Injection and Restricted Pseudo-Labeling
After pre-training and episodic fine-tuning are completed and the semantic injection network has been trained, we evaluate the model on $N$-way $K$-shot test episodes composed of a labeled support set $\mathcal{S}_{\text{test}}$, an unlabeled pool $\mathcal{U}_{\text{test}}$, and a query set $\mathcal{Q}_{\text{test}}$. We first run CVOC on $\mathcal{S}_{\text{test}}\!\cup\!\mathcal{U}_{\text{test}}$ to obtain the visual centers $P(y)$ for each class. For each class $y$ we encode its textual description with the CLIP text encoder and obtain a semantic vector $t_{y}$ [radford2021learning]. We process the concatenated $P(y)$ and $t_{y}$ through the encoder $E_{s}$ of the semantic-injection network yields the refined prototype $P^{\text{ref}}(y)$.
@@
where $s\in[0,1]$ is a fixed blend weight. This constitutes the overall procedure of our semantic anchor module — aligning visual prototypes via the semantic injection network and prototype correction according to Eq. 19. Visual features alone are often ambiguous in few-shot settings. The text embedding acts as a noise-free prior that anchors the visual prototype.
For each unlabeled sample $x_{i}^{u}\in\mathcal{U}_{\text{test}}$ we compute the entropy $H(x_{i}^{u})$ of its softmax logits and retain only the lowest-entropy $k\%$ samples. The selected subset and its hard pseudo-labels are computed as
@@
The support set is then expanded with these high-confidence examples: ${S}^{0}=\mathcal{S}_{\text{test}}\cup\widetilde{\mathcal{U}}_{\text{test}}$.
Balanced label propagation [iscen2019label] is applied to the features of $\mathcal{S}^{0}\cup\mathcal{Q}_{\text{test}}$, using the labels in $\mathcal{S}^{0}$ as anchors. The propagated soft labels for $\mathcal{Q}_{\text{test}}$ are finally hardened by selecting the highest-probability class, yielding the episode predictions (Fig. 3).
@@
Method Venue 1-shot 5-shot 1-shot 5-shot
[…]
TPN ICLR’19 59.46% 75.65% 71.41% 81.12%
[…]
Table 2: Classification accuracy (%) on the tieredImageNet dataset.
@@
Method Venue 1-shot 5-shot 1-shot 5-shot
[…]
PLML-EP IEEE TIP’24 79.62% 86.69%
ICI CVPR’20 85.44% 89.12%
[…]
@@ 5.1 Results
We conducted experiments under 5-way 1-shot and 5-way 5-shot settings, on the miniImageNet, tieredImageNet, and CUB-200-2011 datasets, using WRN-28-10 and ResNet-12 backbones. For the miniImageNet dataset (Table 1), our approach outperformed cluster-FSL [ling2022semi], achieving a 6.7% and 1.4% improvement in 1-shot and 5-shot settings with ResNet-12, and a 3.31% and 1.87% improvement with WRN-28-10. On the tieredImageNet dataset (Table 2), the proposed approach outperforms the closest method by margins of 1.66% and 1.19% for 1-shot and 5-shot cases with ResNet-12, and by a margin of 3.32% for 1-shot and 1.35% for 5-shot settings with WRN-28-10. On the CUB-200-2011 dataset (Table 3), our approach outperforms EPNet [rodriguez2020embedding] and cluster-FSL [ling2022semi] by significant margins. All experiments were conducted on 1000 few-shot tasks with 100 unlabeled data, and 95% confidence intervals are reported.
@@
Method Venue 1-shot 5-shot 1-shot 5-shot
[…]
@@
[…]
@@ 5.2 Ablation Experiments
We experimentally analyze the contributions of components of our method on the miniImageNet dataset and ResNet12 architecture. The results in Table 5 indicate that removing the semantic injection, or intra-inter class distance, significantly degrades the performance of our method. Fig. 4 indicates the utility of CST. Please refer to the Appendix for other ablation experiments and results.
@@ 6 Conclusion
In this work, we proposed a novel SSFSL approach that performs improved clustering-based pseudo-labeling that promotes better intra/inter-class alignment and leverages semantic anchoring. Extensive experiments demonstrate that our method significantly outperforms recent state-of-the-art approaches.
7 Acknowledgements
Assistance through the SERB project SRG/2023/002106 is duly acknowledged.
8 Appendix
8.1 Generalization and Robustness Evaluation
To rigorously evaluate the robustness and generalization capabilities of our proposed method, we conducted a series of experiments under challenging conditions that simulate real-world scenarios. These include performance evaluation in open-set settings with distractor classes and cross-domain few-shot learning tasks. The results underscore the model’s ability to maintain high performance beyond standard benchmark conditions.
8.1.1 Performance in Distractive Settings
A more realistic semi-supervised few-shot learning scenario is the distractive semi-supervised few-shot learning setting, where the unlabeled data includes samples from distractor classes that are different from the classes in the episode. Most existing few-shot learning studies do not consider this setting. We evaluated our approach in this setting on the miniImageNet dataset.
As shown in Table 6, our method achieves a new state-of-the-art performance, outperforming previous approaches. This demonstrates the strong open-set recognition capability of our framework, highlighting its accuracy and stability when faced with unknown classes. For this setting, we follow the experimental setup of [lazarou2025exploiting]. The collection of unlabeled data is broadened by incorporating examples from three additional distractor classes. These classes are distinct and do not overlap with the actual classes in the episode. For the distractor classes, 30 examples are added per class in 1-shot epsiodes/tasks and 50 per class in 5-shot episodes/tasks under the 30/50 configuration. This expansion results in a total of 240 unlabeled examples for 1-shot tasks and 400 for 5-shot tasks in this specific setup. The closest baseline iLPC [lazarou2025exploiting] proposes a synergistic approach of predicting pseudo-labels by leveraging the data manifold and interpreting confident pseudo-label prediction as a form of label cleaning; whereas our approach advances clustering quality, representation learning, and semantic information injection for more accurate pseudo-labeling rather than relying on direct label cleaning from manifold-based confidence.
Table 6: Performance on miniImageNet under the Distractive Semi-Supervised Few-Shot Learning Settings.
Method Network 5-way 1-shot (%) 5-way 5-shot (%)
Masked soft k-means  [ren2018meta] ResNet-12 61.00 72.00
TPN  [liu2018learning] ResNet-12 61.30 72.40
LST  [sung2018learning] ResNet-12 64.10 77.40
MCT  [kye2020meta] ResNet-12 69.60 $\pm$ 0.70 81.30 $\pm$ 0.50
iLPC  [lazarou2025exploiting] WRN-28-10 76.22 $\pm$ 0.86 87.15 $\pm$ 0.43
Ours ResNet-12 74.92 $\pm$ 0.71 85.65 $\pm$ 0.44
Ours WRN-28-10 78.59 $\pm$ 0.66 89.98 $\pm$ 0.36
8.1.2 Generalization to Domain Shift
To assess the ability of our approach to generalize to new data distributions, we performed a stringent cross-domain evaluation. The ResNet-12 backbone, pre-trained and fine-tuned only on miniImageNet, was tested on several target datasets with significant domain shifts: CUB, EuroSAT, ISIC, and ChestX. For these experiments, we evaluate performance with the semantic anchor module disabled (to isolate the core visual framework), as well as enabled with training on either the source domain or a subset of the target data.
The results, presented in Table 7, demonstrate strong cross-domain generalization. Our approach coupled with semi-supervised framework is even competitive with and often surpasses methods specifically designed for domain adaptation eg. self-supervised learning (SSL) methods, even without any training on the target domains. This highlights the intrinsic generalization power of our proposed clustering. We follow [oh2022understanding] to rank our target datasets. The established order for domain similarity to ImageNet is CUB > EuroSAT > ISIC > ChestX, and for few-shot difficulty is ChestX > ISIC > CUB > EuroSAT.
Based on the results, we can deduce a particular failure mode of the semantic anchor module: the module fails to achieve optimal performance when relying solely on source-domain (miniImageNet) training, requiring access to in-domain target data to fully resolve geometric misalignment and generate the highest results.
Table 7: Cross-Domain Few-Shot Classification Accuracy (5-way 1-shot). We report the results for the SSL methods from [oh2022understanding].In contrast to SSFSL setup, transductive FSL setup has no additional unlabeled set. SA: Semantic Anchor.
Method Network CUB EuroSAT ISIC ChestX
SSL pre-training on target domain
SimCLR [Chen2020SimCLR] ResNet-18 - 84.30% 36.39% 21.55%
MoCo [He_2020_CVPR] ResNet-18 - 69.11% 29.54% 21.74%
BYOL [Grill2020Bootstrap] ResNet-18 - 66.16% 34.53% 22.75%
SimSiam [Chen_2021_CVPR] ResNet-18 - 70.80% 30.17% 22.17%
No target domain training; trained on miniImageNet; Transductive FSL inference
RDC [Li_2022_CVPR] ResNet-10 47.77% 67.58% 32.29% 22.66%
LDP-net [Zhou_2023_CVPR] ResNet-10 55.94% 73.25% 33.44% 22.21%
No target domain training; trained on miniImageNet; SSFSL inference; without SIN
Cluster-FSL [ling2022semi] ResNet-12 56.83% 70.78% 33.72% 21.81%
Ours ResNet-12 60.95% 75.66% 36.25% 21.90%
No target domain training; trained on miniImageNet; SSFSL inference; with SA trained on miniImageNet
Ours ResNet-12 61.45% 75.86% 36.20% 21.73%
No target domain training; trained on miniImageNet; SSFSL inference; with SA trained on subset of target data not in test target data
Ours ResNet-12 62.75% 77.36% 37.10% 22.42%
8.1.3 Qualitative Analysis via Class Activation Mapping
To offer a qualitative insight into the generalization of the learned representations for our approach, we visualize and compare Class Activation Maps (CAMs) on a target domain unseen during training. Specifically, we evaluate models that were pre-trained and fine-tuned exclusively on miniImageNet, and then generate CAMs using Grad-CAM as proposed by Selvaraju et al. [selvaraju2017grad] on images from the CUB dataset. This cross-domain setup provides a stringent test of how well the learned feature space transfers to novel data distributions. We selected Cluster-FSL by Ling et al. [ling2022semi] as our primary baseline due to its methodological similarity since it also employs a clustering-based strategy in the finetuning stage. This makes it an ideal candidate for comparing the generalization quality of our proposed Class-Variance Optimized Clustering (CVOC).
As shown in Figure 5, our approach demonstrates superior cross-domain generalization. On the out-of-domain CUB dataset, its activations are more semantically focused on the target object, unlike the diffuse activations from Cluster-FSL. This indicates that our method promotes a more robust and transferable representation, which directly facilitates more accurate clustering on novel domains and underpins our strong quantitative results.
Original Image
Cluster-FSL
Ours
Figure 5: Qualitative comparison of cross-domain activation heatmaps on the CUB dataset. All models were trained on miniImageNet. The columns show the original image, the heatmap from Cluster-FSL (often diffuse), and the heatmap from our approach (superior localization), respectively.
8.2 Notation Table
The notations used throughout this paper are summarized in Table 8.
Table 8: List of notations used in this paper.
Symbol Description
$C_{b},C_{f}$ Base classes and novel/few-shot classes.
$D_{tr},D_{va},D_{tt}$ Training, validation, and test datasets.
NN Number of classes in an episode (“N-way”).
$K$ Number of support samples per class (“K-shot”).
$S=\{(x_{i}^{s},y_{i}^{s})\}_{i=1}^{N\times K}$ Support set (labeled examples).
$Q=\{(x_{i}^{q},y_{i}^{q})\}_{i=1}^{N\times T}$ Query set (test samples).
$U=\{x_{i}^{u}\}_{i=1}^{N\times u}$ Unlabeled samples in an episode.
$M$ Visual feature extractor network.
$S$ (network) Semantic injection network with encoder $E_{s}$ and decoder $D_{s}$.
$s^{(c)}_{k}$ $k$-th support embedding of class $c$.
$P^{(c)}_{0}$ Initial prototype of class $c$ (mean of its support embeddings).
$P(c)$ Current prototype of class $c$ (updated during clustering).
$\beta^{*}_{c}(x)$ Optimal reconstruction weights of feature $x$ with respect to class $c$.
$d_{\text{rec}}(x,c)$ Reconstruction distance of feature $x$ w.r.t. class $c$.
$ep(M(x))$ Embedding-propagated feature of $x$.
$L_{\text{intra}}$ Intra-class distance loss.
$L_{\text{inter}}$ Inter-class distance loss.
$d(x,c)$ Combined distance (reconstruction + intra/inter regularization).
$w_{\text{intra}},w_{\text{inter}}$ Weights for intra-class and inter-class regularizers.
$\epsilon$ Small positive constant (numerical stability, prototype margin).
$\tau$ Softmax temperature parameter.
$B(c)$ Brightness score of prototype $c$ in the Cluster Separation Tuner (CST).
$\beta,\beta_{0},\gamma,\alpha$ CST parameters: attractiveness, maximum attraction, spatial decay, and random exploration amplitude.
$L_{cvoc}$ Cross-entropy loss with CVOC logits.
$L_{lp}$ Cross-entropy loss with label propagation logits.
$L_{cls}$ Cross-entropy classification loss on all embeddings.
$w_{cls},w_{fs}$ Loss weights in episodic objective.
$\eta$ Weight balancing CVOC and label propagation in episodic loss.
$L_{episode}$ Final episodic loss combining all terms.
$L_{s}$ Total loss for training semantic injection network.
$L^{jep}_{s},L^{fr}_{s},L^{sr}_{s}$ Semantic injection losses: joint embedding prototype, feature reconstruction, and semantic reconstruction.
$P^{r}_{c},s^{r}_{c}$ Reconstructed prototype and semantic feature for class $c$.
$P^{ref}(y)$ Refined prototype of class $y$ via semantic injection.
$s$ Mixing weight between visual prototype and semantic embedding in semantic injection.
$H(x)$ Entropy of logits for unlabeled sample $x$, used in restricted pseudo-labeling.
8.3 Experimental Details
8.3.1 Datasets